# gamecenter,boxed

### v1

@@ 5.2.4. Global Update: Masked Secure Aggregation
Figure 2. Global Masked Secure Aggregation Protocol with $M$ Clients.
Server Aggregates Data
@@
Upon receiving the masked matrices, the server first sums over each client’s contribution in Line 2. This yields the masked global sum $\boldsymbol{v}+\boldsymbol{r{}}$. The server then samples from a Laplace distribution to add noise proportional to $\Delta/\epsilon$ for each entry in the result matrix. This noise must also be converted to fixed-point by computing $\tilde{\gamma}=\lceil{\gamma\times\mathbf{2^{\delta}}}\rfloor$. We show why this preserves differential privacy in Appendix A. Finally, the server adds the noise $\tilde{\gamma}$ to the sum and broadcasts it back to the clients.
Client Unmasks Data
@@
Upon receiving the result back from the server, each client must unmask the result. They do this by simply subtracting the mask $r{}$ that they computed in Line 2. After unmasking, each client must scale down the result to retrieve the correct answer. This is done by dividing the unmasked sum by the scale factor, $\mathbf{2^{\delta}}$, to reverse the initial scaling operation.
5.2.5. Postprocessing
In Algorithm 1, each client locally postprocesses the results (since adding noise to the centroids can result in values outside of the domain). A naive postprocessing strategy is to simply truncate out-of-bounds centroids to the boundary. However, in practice, we find that folding (Holohan et al., 2019) the value (reflecting it over the boundary) gives better utility. More formally, if $x<-B$, we adjust it to $x^{\prime}=-2B-x$, and if $x>B$, we adjust it to $x^{\prime}=2B-x$. This folding process effectively “folds” (mirrors) the value over the nearest boundary, preserving its distance from the boundary. We show in Appendix B that folding improves the utility of the algorithm, especially when the noise is large (small privacy budget).
5.3. Proof of Privacy
@@
FastLloyd allows for a central party to add noise, retaining the utility of the central model of DP. However, since the server only interacts with masked values, and the clients can only observe a differentially private view of the final (noised) centroids, the protocol provides a level of privacy akin to that of the local model of DP. We prove the end-to-end privacy of our algorithm in the following theorem.
@@
Algorithm 1 ensures $\epsilon(\lambda)$-IND-CDP-MPC in the presence of a semi-honest, polynomial time adversary who controls at most a single party. Where
$\epsilon=Td\epsilon^{\mu}$
Proof.
In order to prove the algorithm satisfies Definition 3.1, we need to consider the view from each party. We begin with the view of the server. First, let us assume that each client samples the random mask $r_{i}$ from a random oracle over $\mathbf{R}$. Under this assumption, the $Enc(\cdot)$ function satisfies information-theoretic security as it is a one-time pad using $r_{i}$ as the pad. In practice, we implement the sampling of $r_{i}$ using a PRNG and thus reduce from information-theoretic to computational security with a negligible term $negl(\lambda)$. Therefore, the view of the server satisfies $\epsilon(\lambda)$-IND-CDP-MPC with $\epsilon=0$.
In regards to the clients, we assume no client colludes with any other client. Then, a client’s view consists of their own dataset $D_{A}$, and the output from each iteration of the protocol. We note that the initialization is data-independent and thus is indistinguishable between neighbouring datasets. Each iteration consists of the assignment and updating of the cluster centroids. The assignment step is simply postprocessing on the previous iteration (or random initialization) and thus incurs no additional privacy cost. The updating of the cluster centroids applies the Laplace mechanism to each dimension separately, with the sensitivity parameter set following Eqn 6 and $\epsilon^{\mu}$ as the privacy budget. Using naive, sequential composition over the dimensions and the iterations, the client’s view is $Td\epsilon^{\mu}$-IND-CDP-MPC, with no computational assumption (as we use the information-theoretic DP properties of the Laplace mechanism and composition (Dwork and Roth, 2014)). Taking the worst case over the clients and the server, the result follows. ∎
5.4. Error Analysis
To analyze the error of our approach, we follow a similar analysis to Su et al. (Su et al., 2016). We analyze the mean-squared error between the true centroid ($\mu_{j}$) and the differentially-private centroid ($\bar{\mu_{j}}$) for one iteration across all dimensions. Substituting the definition of the centroid method (Eqn 7), we get
$\mathsf{MSE}\left(\bar{\mu_{j}}\right)$ $=\mathsf{E}\left[\sum_{h=1}^{d}((\mu_{j}[h]+\gamma^{\mu})-\mu_{j}[h])^{2}\right]=$
$=\sum_{h=1}^{d}\mathsf{E}\left[(\gamma^{\mu})^{2}\right]=d\cdot\mathsf{Var}\left(\gamma^{\mu}\right).$
Where the last line holds because $\gamma^{\mu}$ is an independent Laplacian variable with zero mean and so
$\mathsf{E}[(\gamma^{\mu})^{2}]=\mathsf{Var}\left(\gamma^{\mu}\right)-(\mathsf{E}[\gamma^{\mu}])^{2}=\mathsf{Var}\left(\gamma^{\mu}\right).$
Plugging in the sensitivity of the mean (Eqn 6) and the privacy budget gives
@@ 5.4. Error Analysis
(9) $\mathsf{MSE}\left(\bar{\mu_{j}}\right)$ $=d\cdot\mathsf{Var}\left(\gamma^{\mu}\right)=\frac{2dB^{2}}{(\epsilon^{\mu}\eta_{min})^{2}}$
(10) $=\frac{2d^{3}T^{2}B^{2}}{\epsilon^{2}\eta_{min}^{2}}=\mathcal{O}\left(\frac{d^{3}T^{2}}{\epsilon^{2}\eta_{min}^{2}}\right).$
This shows that the distortion in the centroid is proportional to $d^{3}T^{2}$ and inversely proportional to $\epsilon^{2}\eta_{min}^{2}$.
Remark.
This error is similar to what Su et al. (Su et al., 2016) observed for the sum-count method with the following differences: (1) $\eta_{min}$replaces the assumed expected cluster size of $\frac{N}{k}$. (2) The error is independent of the data or the actual clustering result; it depends completely on the parameters set before running the algorithm. (3) This is a worst-case analysis, and the actual error might be lower.
Adding across all clusters gives
(11) $\mathsf{MSE}\left(\bar{\boldsymbol{\mu}{}}\right)=\frac{2kd^{3}T^{2}B^{2}}{\epsilon^{2}\eta_{min}^{2}}.$
Using this analysis, we can determine the optimal number of iterations, following a similar approach as Su et al. (Su et al., 2016). First, we calculate a minimum privacy budget per iteration ($\epsilon^{(m)}$) so that the mean-squared error is within a certain bound. Secondly, we divide the available budget ($\epsilon$) by $\epsilon^{(m)}$ in order to determine the number of iterations $T$. Thus, setting $T=1$ and $\epsilon=\epsilon^{(m)}$ gives us
$\mathsf{MSE}\left(\bar{\boldsymbol{\mu}{}}\right)=\frac{2kd^{3}B^{2}}{(\epsilon^{(m)}\eta_{min})^{2}}\leq 0.004$
which implies
$\sqrt{\frac{500kd^{3}B^{2}}{\eta_{min}^{2}}}\leq\epsilon^{(m)}.$
@@
However, Su et al. note that, in practice, there is no need to go beyond seven iterations and that at least two iterations are needed to gain useful results. Therefore, we truncate the number of iterations to be in the range $[2,7]$.
6. Evaluation
@@
    How does the utility of FastLloyd  compare with state-of-the-art in central model (Su et al., 2016) for DP $k$-means?
@@
    How does constrained clustering affect the utility of the sum-count method?
@@
    How does the runtime, communication, and number of rounds for FastLloyd compare with the state-of-the-art in federated $k$-means using secure computation (Mohassel et al., 2020)?
  5. Q5
@@
6.1. Experimental Setup
(a) S1
(b) Lsun
(c) Birch2
(d) House
(e) Iris
(f) Adult
Figure 3. Utility evaluation of FastLloyd vs. SuLloyd on real datasets (20 runs).
Implementaion
The experiments were carried out on a Macbook Pro M2 Max, with a 30-core CPU, a 38-core GPU, and 64 GB of RAM. We use the Open MPI (Gabriel et al., 2004) (Open Source High-Performance Computing) library to implement multiparty communication through message-passing. All experiments were conducted in the multiparty framework, with datasets being randomly split between parties.
Following Mohassel et al. (Mohassel et al., 2020), we only evaluate in the LAN setting, and observe that WAN runtimes can be computed with the linear cost model (adding a fixed latency for each round). We simulate the network latency by introducing a custom (sleep) delay of 0.5ms for every “send” operation. We make our code publicly available to reproduce the results222 https://github.com/D-Diaa/privateclustering.git. All results were averaged over 20 runs and the 95% mean confidence interval is supplied where appropriate.
Baselines
To evaluate the clustering quality of FastLloyd, we compare it against the state-of-the-art in central DP for $k$-means clustering by Su et al.  (Su et al., 2016), denoted as SuLloyd. As established in Section 4, SuLloyd is the most accurate DP $k$-means clustering algorithm to date. We implement a federated version of SuLloyd in our framework for fair evaluation. To study the effect of constrained clustering on the sum-count method used in SuLloyd, we also extend the federated implementation of SuLloyd to use the same constraints as FastLloyd, denoted as CSuLloyd. Finally, we give a non-private baseline of the original Lloyd’s algorithm (Lloyd, 1982), implemented into our federated framework (with sphere packing initialization) and denoted as Lloyd.
To evaluate the computational and communicational costs in the secure federated setting, we compare against the corresponding state-of-the-art by Mohassel et. al (Mohassel et al., 2020), denoted as MohLloyd. According to the SoK of Hedge et al. (Hegde et al., [n. d.]), MohLloyd is the most efficient secure $k$-means clustering algorithm to date. For MohLloyd, we evaluate against the reported values in the original paper (Mohassel et al., 2020).
Parameters
In our experiments, we use the following default parameters, unless otherwise specified. The number of clients is set to two, and the privacy budget $\epsilon=1$. For the MSA protocol (Figure 2), the underlying ring $\mathbf{R}$ is $\mathcal{Z}_{2^{32}}$ and the values are all transformed to fixed-point under precision factor $\delta=16$. For any dataset of size $N{}$, to be clustered into $k{}$ clusters, we fix $\eta_{min}=\frac{N{}}{1.25k{}}$ and $\eta_{max}=\frac{1.25N{}}{k{}}$ (empirically chosen as per Appendix B).
Datasets
We selected a wide range of commonly used datasets to evaluate our approach, we follow Su et al. (Su et al., 2016) and Mohassel et al. (Mohassel et al., 2020) to maximize comparability. An exception is the Birch2 (Zhang et al., 1997) dataset, which is a synthetic dataset comprised of 100 Gaussian clusters forming a sine wave. This dataset initially contains 100K data points, however, a random sample of 25K points is used. To the best of our knowledge, we are the first federated privacy-preserving work to evaluate using a dataset of this size (in terms of $N$ and $k$).
Additionally, we created synthetic datasets to evaluate the scalability of our approach. Similar to Su et al. (Su et al., 2016), we use the R package (Qiu and Joe, 2023), clusterGeneration. The package is designed for generating cluster datasets with a given degree of separation ($\in[-1,1]$) between neighbouring clusters. For the clustering quality evaluation, we generate two sets of datasets: one with overlapping cluster structure (SynthOV, separation = 0.21) and one with well-separated cluster structure (SynthSEP, separation = 0.70). Each set varies $k$ and $d$ from $2$ to $20$ in steps of $2$, resulting in 100 variations. For every variation, we generate three different datasets with different seeds. For all synthetic datasets, we fix an average cluster size of $C_{avg}=\frac{10K}{k}$ and randomly sample cluster sizes to be in the range $[0.75\cdot C_{avg},1.25\cdot C_{avg}]$, resulting in a dataset that has a total size in the range $[7.5K,12.5K]$. For our timing evaluation, we generate similar synthetic datasets, TimeSynth, similar to those in Mohassel et al.’s work (Mohassel et al., 2020), enforcing an exact cluster size of $C_{avg}=\frac{N}{k}$. We vary $N$ between 10K and 100K, $d$ between two and five and $k$ between two and five. A summary of these datasets is presented in Table 2. As is common in the DP literature (Blum et al., 2005; McSherry, 2009; Dwork, 2011; Su et al., 2016), we assume all datasets are normalized in the range $[-1,1]$ as pre-processing.
Table 2. Summary of Datasets Used in Evaluation

### v2

@@ 5.2.4 Global Update: Masked Secure Aggregation
Figure 1: Global Masked Secure Aggregation Protocol with $M$ Clients.
@@
Server Aggregates Data. Upon receiving the masked matrices, the server first sums over each client’s contribution in Line 1. This yields the masked global sum $\bm{v}+\bm{r{}}$. The server then samples from a zero mean Gaussian distribution with standard deviation equal to the supplied $\sigma$ for each entry in the result matrix. This noise must also be converted to fixed-point by computing $\tilde{\gamma}=\lceil{\gamma\times 2^{q}}\rfloor$. We show why this preserves differential privacy in Appendix A.3. Finally, the server adds the noise $\tilde{\gamma}$ to the sum and broadcasts it to the clients.
@@
Client Unmasks Data. Upon receiving the result from the server, each client must unmask the result. They do this by simply subtracting the mask $r{}$ that they computed in Line 1. After unmasking, each client must scale down the result to retrieve the correct answer. This is done by dividing the unmasked sum by the scale factor, $2^{q}$, to reverse the initial scaling operation.
5.2.5 Post-processing
In Algorithm 1, each client locally post-processes the results to obtain the centroids for the next iteration.
First, each party divides the relative sum by the count and shifts the result by the previous centroid to get the new centroid (Line 19) following Eqn 10. Then, if the new centroid is more than $\eta$ away from the previous centroid, we truncate it to be $\eta$ away.
Finally, we apply a post-processing step to the centroids to ensure they remain within the domain. A naive post-processing strategy is to simply truncate out-of-bounds centroids to the boundary. However, in practice, we find (in Appendix B) that folding [32] the value (reflecting it over the boundary) gives better utility. More formally, the operation folds a value $x$ into the range $[-B,B]$ through modular arithmetic. We first compute $(x+B)\bmod(2B)$, then if this result exceeds $B$, we reflect it about $B$ by subtracting it from $2B$. Finally, we return the value to the target range by subtracting $B$. This approach creates a periodic folding pattern that naturally reflects values across the boundaries while preserving distances from the nearest boundary point. We show in Appendix B that folding improves the algorithm’s utility.
5.3 Privacy Analysis
@@
FastLloyd allows for a central party to add noise, retaining the utility of the central model of DP. However, since the server only interacts with masked values, and the clients can only observe a differentially private view of the final (noised) centroids, the protocol provides a level of privacy akin to that of the local model of DP.
We state the end-to-end security Theorem of our algorithm and defer the proof to Appendix A.2.
@@
Algorithm 1 ensures $(\epsilon(\lambda),\delta)$-IND-CDP-MPC in the presence of a semi-honest, polynomial time adversary who controls at most a single party.
5.4 Error Analysis
To analyze the error of our approach, we follow a similar approximate error analysis as Su et al. [64]. The purpose of the analysis is primarily to choose the ratio of the privacy parameters and the number of iterations. Thus, the analysis makes a series of approximations. Following Su et al., we consider a single iteration and a single cluster for this analysis. To simplify the notation, we omit the cluster index $j$ and instead index the variables by the dimension $h$. We analyze the mean-squared error between the true centroid ($\mu$) and the differentially-private centroid ($\tilde{\mu}$) for one iteration across all dimensions.
$MSE(\tilde{\mu}^{(t)})=\mathbb{E}\left[\sum\limits_{h=1}^{d}(\mu_{h}^{(t)}-\tilde{\mu}_{h}^{(t)})^{2}\right]$ (15)
We first expand the following term using the definition of our DP mechanism from Section 5.1:
$\mu_{h}^{(t)}-\tilde{\mu}_{h}^{(t)}$ $=$ $\frac{(C^{(t)}+\gamma^{C})\mu_{h}^{(t)}}{C^{(t)}+\gamma^{C}}-\frac{S_{h}^{(t)}+\gamma^{R}_{h}+\gamma^{C}\tilde{\mu}_{h}^{(t-1)}}{C^{(t)}+\gamma^{C}}$
$=$ $\frac{\gamma^{C}(\mu_{h}^{(t)}-\tilde{\mu}_{h}^{(t-1)})-\gamma^{R}_{h}}{C^{(t)}+\gamma^{C}}.$
Let $\nabla_{h}$ = $\mu_{j}^{(t)}-\tilde{\mu}_{j}^{(t-1)}$. Then, after squaring, we get:
$(\mu_{h}^{(t)}-\tilde{\mu}_{h}^{(t)})^{2}=\frac{(\gamma^{C})^{2}(\nabla_{h})^{2}-2\gamma^{C}\gamma^{R}_{h}\nabla_{h}+(\gamma^{R}_{h})^{2}}{(C^{(t)}+\gamma^{C})^{2}}.$
After taking the expectation over all clusters and assuming that $C+\gamma^{C}\approx N/k$, following Su et al. [64], we get:
@@ 5.4 Error Analysis
$\mathsf{MSE}\left(\tilde{\mu}_{h}^{(t)}\right)$$\approx$ $\frac{k^{3}}{N^{2}}\left(\mathbb{E}\left[(\gamma^{C})^{2}\right]\mathbb{E}\left[\nabla_{h}\right]^{2}+\mathbb{E}\left[(\gamma^{R}_{h})^{2}\right]\right)$
$=$ $\frac{k^{3}}{N^{2}}\left(\mathsf{Var}\left(\gamma^{C}\right)\mathbb{E}\left[\nabla_{h}\right]^{2}+\mathsf{Var}\left(\gamma^{R}_{h}\right)\right)$
where, in the first line, the middle term goes is zero as the noise terms are i.i.d. and zero mean. The second line holds because each $\gamma$ is an independent variable with zero mean, and so
$\mathsf{E}[(\gamma)^{2}]=\mathsf{Var}\left(\gamma\right)-(\mathsf{E}[\gamma])^{2}=\mathsf{Var}\left(\gamma\right).$
Finally, we approximate $\mathsf{E}[\sum_{h=1}^{d}\nabla_{h}]\approx\frac{\eta}{2}$. We argue that if the data were uniformly distributed in a hypersphere of radius $\eta$ around the centroid, the expected distance from the centroid would be $\eta/2$. This gives the final error term over all dimensions:
$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}}{4N^{2}}\left(\mathsf{Var}\left(\gamma^{C}\right)\eta^{2}+4d\mathsf{Var}\left(\gamma^{R}\right)\right)$ (16)
Privacy Budget Splitting
We now use the approximate analysis in Eqn 16 to determine the optimal privacy budget split between the relative sum and the count. We substitute the variance of the noise terms from Eqn 7 into Eqn 16:
$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)\approx\frac{k^{3}\eta^{2}T}{4N^{2}}\left((\sigma^{C})^{2}+4d(\sigma^{R})^{2}\right).$ (17)
In our privacy analysis, we compute the noise multiplier $\sigma$ such that we achieve $\frac{1}{\sigma}$-GDP. Thus, we need to split the noise multiplier between the relative sum and the count following Theorem 2.2:
$\sqrt{\left(\frac{1}{\sigma^{R}}\right)^{2}+\left(\frac{1}{\sigma^{C}}\right)^{2}}=\frac{1}{\sigma}$ (18)
We minimize the per iteration error, subject to this constraint, using Lagrange multipliers, which gives:
$\sigma^{C}=\sqrt[4]{4d}\sigma^{R}.$ (19)
Simply scaling each sigma by this ratio gives Eqn 7. Substituting Eqn 7 back into the error analysis gives:
$\mathsf{MSE}\left(\tilde{\mu}^{(t)}\right)=\frac{k^{3}\eta^{2}T\sigma^{2}(1+\sqrt{4d})^{2}}{4N^{2}}$ (20)
Optimal Number of Iterations
Using this analysis, we can determine a heuristic for the number of iterations. Following Su et al. [64], we assume that the error in each iteration is less than $0.004\times B$. Re-arranging for $T$ gives:
$T<\frac{4N^{2}(0.004)}{k^{3}\eta^{2}\sigma^{2}(1+\sqrt{4d})^{2}}.$ (21)
@@
However, as Su et al. [64] explain, in practice, there is no need to go beyond seven iterations and at least two iterations are needed to gain useful results. Therefore, we truncate this analysis such that the number of iterations is in the range $[2,7]$.
5.5 Setting the Maximum Radius Parameter
We derive a dimensionality-aware radius parameter $\eta$ based on the geometric properties of the feature space. Consider a $d{}$-dimensional bounded feature space where each dimension is constrained to the interval $[-B{},B{}]$. The domain diagonal $\beta$, representing the maximum possible distance between any two points in this space, is given by:
$\beta=\sqrt{d{}}\times(2B{})$ (22)
When partitioning this space into $k{}$ clusters, each cluster occupies a fraction of the total volume. The total volume of the feature space is $(2B{})^{d{}}$. Assuming uniform partitioning, each cluster occupies a volume of $(2B{})^{d{}}/k{}$. Consequently, the effective length of each cluster along any dimension can be expressed as:
$L_{c}=\frac{2B{}}{k{}^{1/d{}}}$ (23)
Assuming hypercube-shaped clusters, the maximum distance from a centroid to any point within the cluster (the cluster radius) is half the cluster’s diagonal. This can be calculated as:
$\text{Cluster Radius}=\frac{L_{c}}{2}\times\sqrt{d{}}=\frac{B{}}{k{}^{1/d{}}}\times\sqrt{d{}}=\frac{\beta}{2k{}^{1/d{}}}$ (24)
Based on this analysis, we propose a heuristic $\eta_{d{}}$ to constrain the cluster radius in a $d{}$-dimensional space as:
$\eta_{d{}}=\frac{\alpha\beta}{2k{}^{1/d{}}}$ (25)
where $\alpha$ is a scaling factor we set to $0.8$ based on experiments using synthetic data in Appendix B.
In scenarios with low $d$ and high $k$, $\eta$ becomes increasingly restrictive as $k{}^{1/d{}}$ grows larger, providing tighter bounds on cluster radii. However, as dimensions increase, the curse of dimensionality necessitates larger cluster radii to accommodate the exponential growth of volume in the space, $(2B{})^{d{}}$. Our approach takes this into account through $k{}^{1/d{}}$ approaching unity in high dimensions. This dimensional scaling aligns with the intuition that radius constraints are most meaningful in lower-dimensional spaces, where cluster boundaries remain well-defined, whereas, in high-dimensional spaces, the curse of dimensionality renders such constraints increasingly less effective as distance metrics lose their discriminative power.
We consider two distinct approaches to implement the radius constraint. In the “Constant” approach, the radius constraint is consistently enforced throughout the clustering process. Alternatively, the “Step” approach initializes with a broader constraint of $\beta/2$ at iteration zero and then transitions to the computed radius $\eta$ for the remaining iterations. This two-phase strategy allows for initial flexibility in centroid placement while gradually imposing stricter constraints for fine-tuning the centroids. We show in Appendix B that the “Step” approach is superior.
6 Evaluation
@@
    How does the utility of FastLloyd  compare with state-of-the-art in central model [64] for interactive DP $k$-means?
@@
    How does the runtime, communication, and number of rounds for FastLloyd compare with the state-of-the-art in federated $k$-means using secure computation [50]?
@@
(a) Adult
(b) Birch2
(c) Iris
(d) Breast
(e) Yeast
(f) House
(g) LSun
(h) S1
(i) Wine
(j) MNIST
Figure 2: Normalized Intra-cluster Variance (NICV) vs $\epsilon$ for real datasets.
6.1 Experimental Setup
6.1.1 Implementation
Our experimental evaluation was conducted on a Macbook Pro M2 Max (30-core CPU, 38-core GPU, 64GB RAM) using Open MPI [26] for multiparty communication. Following Mohassel et al. [50], we evaluate in the LAN setting with simulated network latency (0.25ms per send operation), noting that WAN runtimes can be derived using a linear cost model. Our experimental framework utilizes a two-client setup with balanced dataset partitioning. The MSA protocol (Figure 1) operates over the ring $\mathbf{R}=\mathcal{Z}_{2^{32}}$, employing fixed-point representation with precision factor $q=16$. All experiments use randomly partitioned datasets, with results averaged over 100 runs and reported with 95% confidence intervals where applicable. For FastLloyd, we set the radius constraint $\alpha=0.8$ using the “Step” strategy (see Appendix B for detailed analysis and justification of these parameter choices). We set $\delta=\frac{1}{N\log{N}}$ for all experiments. Our implementation is publicly available at https://github.com/D-Diaa/FastLloyd.
6.1.2 Baselines
We evaluate FastLloyd against several baselines. For evaluating utility, we implement three protocols. First, Lloyd [44], a non-private Lloyd’s algorithm adapted to our federated framework with sphere packing initialization. Second, SuLloyd [64], which adapts Su et al.’s centralized differentially private $k$-means clustering algorithm to our federated setting using our masked secure aggregation (MSA). Third, GLloyd, a modification of SuLloyd replacing Laplace noise with Gaussian noise and using composition and privacy budget analysis similar to our work. In Appendix B.1, we provide the details of how the analysis in Section 5.4 differs for GLloyd. For computational and communication efficiency benchmarking, we compare against MohLloyd [50], established by Hedge et al. [31] ass the most efficient secure $k$-means protocol.
6.1.3 Datasets
Our evaluation employs both real and synthetic datasets. We use established datasets from the clustering datasets repository [25], following Su et al. [64] and Mohassel et al. [50] for comparability. For Birch2 [72], we take $25,000$ random samples from the $100,000$ sample dataset.
For evaluating scalability, following [64], we generate synthetic datasets (Synth) using the clusterGeneration R package [58], which enables control over inter-cluster separation (in $[-1,1]$). The synthetic datasets contain $N=10,000$ samples across $k$ clusters, with cluster sizes following a $1:2:...:k$ ratio. We incorporate a random number (in $[0,100]$) of randomly sampled outliers and set the cluster separation degrees in $[0.16,0.26]$, spanning partially overlapping to separated clusters. While Su et al. [64] evaluate configurations up to $k=10$ and $d=10$, we extend the evaluation to $k=32$ and $d=512$ in powers of two, creating $45$ parameter combinations. Each combination generates three datasets with different random seeds. To assess scalability at higher cluster counts, we further extend our evaluation with Synth-K, incorporating configurations up to $k=128$ for $d=2$.
For benchmarking performance, we create synthetic datasets (TimeSynth) with balanced cluster sizes ($C_{avg}=\frac{N}{k}$), varying $N$ (10K-100K), $d$ (2-5), and $k$ (2-5), following Mohassel et al. [50]. Table 2 summarizes these datasets. Following standard practice in DP literature [9, 47, 16, 64], all datasets are normalized to $[-1,1]$.
Table 2: Summary of Datasets Used in Evaluation